# QuEST: Low-bit Diffusion Model Quantization via Efficient Selective Finetuning

### v3

@@ 3.2 Experiment Results and Analysis
Table 7 compares efficiency and performance between our method and the post-training quantization approach on the LSUN-Bedrooms dataset when using LDM-4. Our approach, while utilizing the same quantity of calibration data as the PTQ method, is more time-efficient and incurs only a 20% increase in GPU memory usage. We also show how each proposed component improves the generation performance. The results indicate that sequentially finetuning time embedding layers followed by finetuning attention-related layers leads to consistent performance improvements.
Table 7: Component and efficiency comparisons on LDM-4 (LSUN-Bedrooms 256 $\times$ 256).
Method
Bit-width
(W/A)
Calibration
data size
Time cost
(GPU hours)
Memory cost
Model size
[…]
FP 32/32 - - - 1045.6 2.95
PTQ [19] 4/8 5120 23.08 10334 148.4 5.37
Baseline 4/8 5120 11.52 9822 148.4 6.95
+ $\text{SLA}_{\text{A}}$ 4/8 5120 13.13 12134 148.4 4.41
+ $\text{SLA}_{\text{A}}$ + $\text{SLA}_{\text{TE}}$ 4/8 5120 15.25 12450 148.4 3.26
Figure 5: Effect of uniformly sampled time steps number. Circle size represents quantization computation cost. Red square indicates the value we choose.
Time step sampling frequency analysis. While time steps are sampled uniformly, will the performance consistently improve when we sample more frequently? Figure 5 illustrates the trade-off between the increased number of sampled time steps (’T-Num’) and generation performance. Metrics are calculated using 10,000 samples generated from LDM-4 on LSUN-Bedrooms. The results show that the though performance improves consistently, the amount of growth gradually decreases. Starting from 5 time steps, using 15 more time steps leads to 25.22 sFID improvement. But using another 30 time steps (from 20 to 50) only leads to 2.92 sFID improvement. Given the linearly increasing time cost associated with using more time steps, we opt for only 20 out of 200 time steps to enhance quantization efficiency.
How QuEST influences model properties. Our approach is based on the certain characteristics observed in quantized diffusion models, hence we aim to analyze how our fine-tuning strategy counters these identified challenges. As shown in Table 2, finetuning time embedding layers leads to even better performance than the full-precision counterpart. We infer that this might be due to both the additional parameters introduced by quantization and the additional task loss supervision. Figure 2 further implies that our method adjusts the activation distribution so that it can be more easily quantized. We see that the ranges of the activation value distributions shrink from [-10, 34] to [-4, 14] and [-11, 20] to [-4, 4]. The standard deviations also decrease from 0.171 to 0.157 and from 0.073 to 0.071. This indicates a higher concentration of values around the mean, which can effectively reduce both rounding and clipping errors.
@@
Integration with LoRA finetuning. Finetuning quantized models is not trivial. Thus, we use QALoRA [11] that extends LoRA [16] and finetune on the ImageNet 256$\times$256 dataset. The weight rank of 32 is used for the LoRA modules and the parameters are trained over 100 time steps for 160 epochs. We find that integrating the QALoRA technique leads to a 5.62 decrease in FID, indicating that finetuning the original layers is a better solution for performance preservation.
@@ 4 Conclusion
We have proposed QuEST, an efficient data-free finetuning framework for low-bit diffusion model quantization. Our method is motivated and guided by the three underlying properties found in quantized diffusion models. We also theoretically demonstrate the sufficiency of finetuning, interpreting it as a way to enhance model robustness against large activation perturbations. To alleviate the performance degradation, we propose to finetune the time embedding layers and the attention-related layers under the supervision of the full-precision counterpart. Experimental results on three high-resolution image generation tasks (including Stable Diffusion) demonstrate the effectiveness and efficiency of QuEST, achieving low-bit compatibility with less time and memory cost.
Appendix A Related Works
A.1 Diffusion Model Inference
Diffusion models [15, 32] generate samples via a learned iterative denoising process. During inference, the initial input is sampled from a Gaussian distribution: $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the final output $x_{0}$ is obtained by removing noises from $x_{T}$ subsequently using:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\tilde{\boldsymbol{\mu}}_{\theta,t}(x_{t}),\tilde{\beta_{t}}\mathbf{I}).$ (9)
In this Markov Chain, $\tilde{\boldsymbol{\mu}}_{\theta,t}$ and $\tilde{\beta_{t}}$ are calculated from the model’s output. For a typical diffusion model, the denoising stage requires hundreds to thousands of iterations. This laborious sequential process makes efficient inference and high-resolution generation extremely challenging. Latent Diffusion Models (LDMs) [32] were developed to overcome the challenges. They encode the data into a latent space and then apply the diffusion process within this space, enabling high-resolution image generation with fewer time steps. The output images are produced by decoding the latent representations. LDM’s concept and network quantization complement each other, mutually enhancing the efficiency of diffusion models.
Practically, diffusion models typically adopt a UNet architecture [33], with some modules substituted by attention blocks. In a latent diffusion model, an encoder and a decoder are incorporated for projecting images into latent representations and subsequently generating images from the latent space, respectively. Usually encoders and decoders are lightweight and computationally inexpensive, so our focus is on quantizing the UNets in LDMs, in alignment with other works (as described below).
A.2 Diffusion Model Quantization
Model quantization is a dominant technique for optimizing the inference memory and speed of deep learning models by reducing the precision of the tensors used in computation. Few researches were done for quantizing diffusion models and they fall into two mainstream frameworks: Quantization-Aware Training (QAT) [17, 20, 21, 38] and Post-Training Quantization (PTQ) [22, 25, 28, 36, 37]. QAT methods such as Q-DM [20] consider value precision as an additional optimization constraint, training all parameters from scratch with the target task objective. While it is effective for low-bit quantization, it is even more resource-intensive than training a full-precision model. Efficient-DM [11] further uses Low-rank adapters (LoRA) [16] to reduce training cost. However, it introduces extra weight parameters and still requires substantial training iterations.
Post-Training Quantization methods aim to calibrate the quantization parameters based on a small calibration set. Pioneering works such as PTQ4DM [34] and Q-Diffusion [19] focus on sampling the outputs of the full-precision model under different time steps with certain probabilities to construct the calibration set. ADP-DM [35] partition time steps into different groups and optimize each group with differentiable search algorithms. PTQ-D [12] decomposes the quantization noise and correct them individually with full-precision model statistics. However, these methods generally fail under 4-bit or lower bit-width. To enable low-bit compatibility with high efficiency, our method adopts the general framework of PTQ but further introduces an efficient weight fine-tuning strategy.
Appendix B Proof for Theorem 2.2
@@ A.2 Diffusion Model Quantization
Since the perturbation $\Delta$ is too large for accurate Taylor expansion, we can resolve it by introducing a new perturbation $\epsilon=\Delta/K$, where we divide $\Delta$ by a constant $K$ so that $\epsilon$ is small enough for approximation. Then, Equation 3 is rewritten as follows:
@@
where the approximation step follows Taylor expansion and only the first two main components are kept. The first term in Equation B cannot be ignored because samples such as $z_{n,t}+(i-1)\epsilon$ may not be included in the learned distribution of the model. The second term can still be minimized by reconstruction since only the difference between quantized model output and ground-truth matters. In the following, we temporarily exclude the second term for simplicity since it can always be minimized through aligning the activation outputs.
@@
where $\mathbf{w}_{n}$ is the weight for layer $n$, $\tilde{z}_{n-1,t}^{i}$ is the activation of the $(n-1)$th layer in a quantized model to get $z_{n,t}+(i-1)\epsilon$. Ground-truth $\overline{z}_{n,t}$ can be approximated by the full-precision output $z_{\text{FP}}$. We see that $\tilde{z}_{n-1,t}^{i}$ and $z_{\text{FP}}$ cannot be changed, thus to minimize Equation B, we need to finetune $\textbf{w}_{n}$. From a general perspective, Equation B also indicates that the model has not converged well to a local minimum given the perturbed inputs, thus when we finetune the model layers given the quantized inputs, we are actually training the model towards convergence over new samples and increasing its robustness.
Appendix C Examples of Imbalanced Activation Distributions
Apart from Figure 2, we show that the imbalance in the activation distribution is a common phenomenon in different model structures and datasets. In Figure 6, we show more results of activation distributions of latent diffusion models on ImageNet 256 $\times$ 256 and LSUN-Bedrooms 256 $\times$ 256.
Appendix D Importance of large values in activations
As shown in Figure 2, quite a few values are rather large and diversely distributed. These values pose difficulties on activation quantization, and being rather important and not negligible. To demonstrate this, we corrupt certain tokens in the activation outputs of the diffusion model and check the corresponding generated images. The corruption is done by setting the token values as all zeros. As shown in Figure 7, we compare two settings: (1) corrupt a certain number of tokens randomly; (2) corrupt the same number of the tokens with the largest values.
Appendix E Generated image examples for different tasks
E.1 Unconditional Image Generation
The generated images for LDM-4 on LSUN-Bedrooms 256$\times$256 under different bit-widths are shown in Figure 8. Images for LSUN-Churches 256$\times$256 are shown in Figure 9.
@@ E.1 Unconditional Image Generation
E.2 Class-conditional image generation
Figure 10 shows the generated images for 3 different classes.
@@ E.2 Class-conditional image generation
Figure 11: Text-to-image generation results on Stable Diffusion.
E.3 Text-to-image generation
Figure 11 shows the generated images for two different prompts using Stable Diffusion v1.4 under different bit-width.
Appendix F Limitations and Broader Impacts

### v4

@@ 4.3 Ablations and Discussions
Efficiency comparison with PTQ methods and the impact of individual components. Tab. 7 compares the efficiency and performance against the post-training quantization (PTQ) approach on the LSUN-Bedrooms dataset. Although our method uses the same amount of calibration data as the PTQ approach, it achieves better time efficiency with only a 20% increase in GPU memory usage. We also illustrate the contribution of each component to generation performance. The results indicate that sequentially finetuning the time embedding layers, followed by attention-related layers, yields consistent performance improvements.
Tab. 7 presents a comparison of performance with and without the global loss $\mathcal{L}_{\text{G}}$. The results indicate that supervising the quantized model using the output difference from the full-precision counterpart is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for TLA and CMA, respectively. However, when the learning process is only supervised by the global loss, we find that the performance degrades by 7.13 FID and 9.39 sFID for TLA, suggesting that the global loss alone is insufficient for optimal performance.
Bit-width Method Time (h) Memory (MB) Iters FID $\downarrow$ EfficientDM 2.60 12004 32k 15.15 Full-finetune 0.85 15076 2.2k 5.38 W4A8 Ours 0.45 12178 2.2k 3.82 EfficientDM 2.60 12004 32k 10.60 Full-finetune 0.85 15076 2.2k 6.36 W4A4 Ours 0.45 12178 2.2k 6.12
Table 8: Efficiency comparison with other finetuning methods.
How QuEST adjusts the activation distribution. Our approach is motivated by the imbalanced activation distribution in diffusion models, hence we aim to analyze how our fine-tuning strategy addresses this challenge. As shown in Fig. 2, our method refines the activation distribution, making it more conducive to quantization. Specifically, the activation value ranges shrink from [-10, 34] to [-4, 14] and from [-11, 20] to [-4, 4]. Additionally, the standard deviations decrease from 0.171 to 0.157 and from 0.073 to 0.071, while the mean remains consistent. This results in a more compact activation distribution, effectively reducing both rounding and clipping errors during quantization.
Comparison with precomputed time embeddings. In diffusion models, time embeddings are independent of input conditions and noise. A potential approach is to precompute these embeddings and reuse them directly. However, this strategy overlooks the compatibility between different modules in a quantized model. We take this into consideration and optimize the time embeddings with $\arg\min_{\mathbf{w}_{l}}(\mathcal{L}_{\text{TLA}}+\mathcal{L}_{\text{G}}),\ \ l\in\mathbb{C}_{\text{TE}}$ so that the time embedding layers are also trained to minimize the final prediction error. As shown in Tab. 2, adding this optimization objective enhances quantization performance, even surpassing the full-precision baseline (which uses precomputed features).
@@
Integration with LoRA finetuning. Different ways exist for finetuning quantized models. We further employ QALoRA [9] to finetune on the ImageNet 256$\times$256 dataset. A rank of 32 is used for the LoRA weights, and the parameters are trained over 100 time steps for 160 epochs. We find that integrating the QALoRA technique leads to a 5.62 increase in FID, indicating that finetuning the original layers is a better solution for performance preservation.
Efficiency comparison with other finetuning methods. We compare with EfficientDM and full-finetuning in terms of actual training costs on LDM-4 in Tab. 8. The setting of full-finetuning is aligned with our method. We observe that: compared with EfficientDM, our method requires fewer training iterations and time to obtain better performance with comparable GPU memory cost. Compared with full-finetuning, our method costs less time and memory, as well as achieving better performance. The bottleneck in computational costs becomes more severe when scaled to larger models such as Stable Diffusion. We find that while full-finetuning quickly encounters OOM, our method is able to finetune SD on a single GPU with 48GB memory.
@@ 5 Conclusion
We have proposed QuEST, an efficient data-free finetuning framework for low-bit diffusion model quantization. Our method is motivated by the current challenge in low-bit diffusion model quantization and guided by the two underlying properties found in quantized diffusion models. To alleviate the performance degradation, we propose to finetune the time embedding layers and the attention-related layers under the supervision of the full-precision counterpart. Experimental results on three high-resolution image generation tasks (including Stable Diffusion) demonstrate the effectiveness and efficiency of QuEST, achieving low-bit compatibility with less time and memory cost.
Supplementary Material
The supplementary material is organized as follows: Sec. 6 provides comparison with TFMQ-DM; Sec. 7 provides comparison on the low-resolution dataset; Sec. 8 provides the proof and detailed analysis for Theorem 3.2; Sec. 9 presents additional examples of the imbalanced distributions across different models; Sec. 10 highlights the importance of the large values in activations; Sec. 11 offers further generated examples from our method across varying bit-widths; and Sec. 12 discusses limitations and broader considerations.
6 More Baseline Comparisons
We further compare with TFMQ [15] below:
Bedroom W8A8 W4A8 TFMQ-DM 3.14 3.68 QuEST 3.03 3.26 ImageNet W8A8 W4A8 TFMQ-DM 10.79 10.29 QuEST 10.43 8.48
Table 9: Comparing TFMQ.
We also supplement the metrics for Table 3:
W8A8 sFID $\downarrow$ IS $\uparrow$ QDiffusion 8.19 2.25 PTQD 9.89 2.25 EfficientDM N/A N/A Ours 6.86 2.27 W4A4 sFID $\downarrow$ IS $\uparrow$ QDiffusion N/A N/A PTQD N/A N/A EfficientDM 15.15 2.27 Ours 7.82 2.26
Table 10: Additional metrics on LSUN-Bedrooms. “N/A” represents generation failure.
7 Low-resolution dataset comparison
We further include experiments on CIFAR10 in Tab. 11.
W8A8 W4A4 Q-Diffusion 3.75 N/A EfficientDM 3.75 10.48 QuEST 3.71 9.37
Table 11: FID comparison on CIFAR10.
8 Proof for Theorem 3.2
@@ 7 Low-resolution dataset comparison
Since the perturbation $\Delta$ is too large for accurate Taylor expansion, we can resolve it by introducing a new perturbation $\epsilon=\Delta/K$, where we divide $\Delta$ by a constant $K$ so that $\epsilon$ is small enough for approximation. Then, Sec. 3.4 is rewritten as follows:
@@
where the approximation step follows Taylor expansion and only the first two main components are kept. The first term in Sec. 8 cannot be ignored because samples such as $z_{n,t}+(i-1)\epsilon$ may not be included in the learned distribution of the model. The second term can still be minimized by reconstruction since only the difference between quantized model output and ground-truth matters. In the following, we temporarily exclude the second term for simplicity since it can always be minimized through aligning the activation outputs.
@@
where $\mathbf{w}_{n}$ is the weight for layer $n$, $\tilde{z}_{n-1,t}^{i}$ is the activation of the $(n-1)$th layer in a quantized model to get $z_{n,t}+(i-1)\epsilon$. Ground-truth $\overline{z}_{n,t}$ can be approximated by the full-precision output $z_{\text{FP}}$. We see that $\tilde{z}_{n-1,t}^{i}$ and $z_{\text{FP}}$ cannot be changed, thus to minimize Sec. 8, we need to finetune $\textbf{w}_{n}$. From a general perspective, Sec. 8 also indicates that the model has not converged well to a local minimum given the perturbed inputs, thus when we finetune the model layers given the quantized inputs, we are actually training the model towards convergence over new samples and increasing its robustness.
9 Examples of Imbalanced Activation Distributions
Apart from Fig. 2, we show that the imbalance in the activation distribution is a common phenomenon in different model structures and datasets. In Fig. 5, we show more results of activation distributions of latent diffusion models on ImageNet 256 $\times$ 256 and LSUN-Bedrooms 256 $\times$ 256.
10 Importance of large values in activations
As shown in Fig. 2, quite a few values are rather large and diversely distributed. These values pose difficulties on activation quantization, and being rather important and not negligible. To demonstrate this, we corrupt certain tokens in the activation outputs of the diffusion model and check the corresponding generated images. The corruption is done by setting the token values as all zeros. As shown in Fig. 6, we compare two settings: (1) corrupt a certain number of tokens randomly; (2) corrupt the same number of the tokens with the largest values.
@@ 10 Importance of large values in activations
Figure 8: Text-to-image generation results on Stable Diffusion.
11 More generated image examples
11.1 Unconditional Image Generation
The generated images for LSUN-Bedrooms 256$\times$256 under different bit-widths are shown in Fig. 7. Images for LSUN-Churches 256$\times$256 are shown in Fig. 9.
@@ 11.1 Unconditional Image Generation
11.2 Class-conditional image generation
Fig. 10 shows the generated images for 3 different classes.
@@ 11.2 Class-conditional image generation
11.3 Text-to-image generation
Fig. 8 shows the generated images using Stable Diffusion v1.4 under different bit-width.
12 Limitations and Broader Impacts